OpenStudy (anonymous):

The sum of two consecutive even integers is 234. What is the larger integer? 114 116 118 124

jimthompson5910 (jim_thompson5910):

hint: solve x+(x+2) = 234 for x

OpenStudy (jdoe0001):

well, if you had an integer of say "a" the consecutive number for that will be "a+1" so the sum of "a" and "a+1" is 234 thus a+(a+1) = 234 just solve for "a" the largest integer will be "a+1" of course :)

jimthompson5910 (jim_thompson5910):

jdoe0001 keep in mind that they want consecutive even integers, not just consecutive integers

OpenStudy (jdoe0001):

ohh ahemm, yeah

OpenStudy (anonymous):

x+(x+2) = 234 2x+2=234 2x=234-2 2x=232 x=232/2 118

OpenStudy (jdoe0001):

right, thus a+(a+2) = 234

jimthompson5910 (jim_thompson5910):

232/2 is not 118 xoya

jimthompson5910 (jim_thompson5910):

you got it jdoe0001